Two persons A and B are rolling a die on the condition that the person who gets 3 will win
the game. If A starts the game, then find the probabilities of A and B respectively to win the
game.

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
P=Probability of getting 3=1/6
q=1-p=5/6
Probability success (p) = 1/6
Probability of failure (q) = 5/6
A may win the game either in I trail or in III trial or in V trial etc.
Probability of A win
= p + q. q. p + q. q. q. q. p + …
=p/(1-q2)=6/11.
∴ Probability of B’s win = 1 – p(A) =5/11
Thanks and Regards
Shaik Aasif
askIITians faculty